# DeepTV: A neural network approach for total variation minimization

### v1

@@ 5.1.3 Evaluation of the error estimate
for all $v\in L^{2}(\Omega)$, where $\rho_{1}(v):=2\|v-g\|_{L^{2}(\Omega)}$ and $\rho_{2}(v):=\frac{1}{\alpha_{2}}\|\xi_{G}\|_{L^{2}(\Omega)}$. In order to investigate the nature of the error estimate and how the two terms $\rho_{1}$ and $\rho_{2}$ behave, we depict $\rho$, $\rho_{1}$ and $\rho_{2}$ together with the real error in figure 8. Note that in figure 8 on the $x$-axes we have the number of updates of the approximation, i.e. the iterations at which the loss was reduced, and not the number of iterations. Exemplary we depict the error estimate for the example in section 5.1.1 with $c=100$ and for the example in section 5.1.2 with $c=10000$ and for $\operatorname{TV}_{2}^{h}$ with Huber-regularization and $\operatorname{TV}_{2,1}^{h}$ with $\gamma$-lifting. A similar behavior can be observed for the other settings of sections 5.1.1 and 5.1.2. In figure 8 we see that while the error estimate indeed always holds, as expected, it is governed by $\rho_{2}$. In fact $\rho_{2}$ leads to an oscillatory behavior of the error estimate $\rho$, which does not seem to decrease although the real error is. This behavior does not really seem to make the error estimate practical for example to be used in a stopping criterion.
@@ 5.2 Applications
We consider three applications, namely denoising, inpainting and deblurring. In all the applications the image is scaled to the unit square $[0,1]^{2}$. In order to recover the original image from the corrupted observation we minimize (19). Thereby in all the experiments we choose a forward difference discretisation of the gradient operator with incorporated homogeneous Neumann boundary conditions. Further we choose $\lambda=1$ and $\operatorname{TV}_{2}^{h}$ with $\gamma$-lifting. That is the last term in (19), i.e. $\frac{\lambda}{N}\sum_{x\in\Omega^{h}}|\nabla^{h}u_{\theta}(x)|_{\blacktriangleright}$, is set to be $\frac{1}{N}\sum_{x\in\Omega^{h}}|\nabla_{\text{F}}^{h}u_{\theta}(x)|_{2,\gamma}$, where $\gamma=10^{-10}$. The solution space $\mathcal{H}_{c}$ consists of ReLU-NN with 3 hidden layers, where each layer has 128 neurons, and we set $c=10000$. Adam with learning rate 0.001 is used to find an approximation and the training is always stopped after 300001 iterations. The parameters $\alpha_{1}$ and $\alpha_{2}$ are selected differently depending on the application and specified at the respective places. However, they are not optimized in any way and only chosen such that the effect of the total variation can be observed. Next, we describe the image reconstruction tasks before we present and discuss the respective results.
@@ Results
The obtained restoration for all applications on the grid of the observed image can be seen in the second column of figures 11 and 12. Since a neural network is a continuous function not restricted to a particular grid, the obtained solutions can be plotted on any discretisation of $\Omega$. We plot the obtained solutions also on a 10 times finer grid, i.e. on a grid with $1290\times 1290$ pixels for the disk-image and $2560\times 2560$ pixels for the building-image, see fourth column of figures 11 and 12. In order to see the difference in representation of the solutions we zoom-in in one part of the images, see third and fifth column of figures 11 and 12. While optimizing on a rather coarse grid, we actually obtain a solution in a continuum allowing us to depict the solution on an arbitrary fine grid. Here we observe that the solution on a fine grid seems to be well adjusted to the nature of the original signal to be restored, e.g. the curvature of the disk seems to be preserved. However, we emphasize that according to propositions 4.7 and 4.8 on a fixed grid the solution of (19) is the same as the respective finite difference solution, if it exists, in the sense that the values of the solutions on the grid points are the same. Hence the effects of the neural network’s design are only apparent or relevant at a level of detail finer than what is captured by the current data resolution.
@@ 6 Conclusion and discussion
We have seen that the considered neural network approach changes the nature of the original problem. In particular, while the $L^{1}$-$L^{2}$-TV model is convex, its neural network version (4) is not convex anymore. Further we observe that even if the original problem has a solution the resulting neural network optimization problem does in general not, cf. example 3.3. Hence, it seems that the neural network approach renders the problem more complicated to deal with. Note that the non-existence of a solution can be easily resolved, as we did by introducing (9). However, the non-convexity seems to be persistent and we are not aware how to circumvent this property. Nevertheless, this additional complexity seems to come with a benefit. First, the solution is mesh-free and continuous allowing to depict the solution on any grid and domain. Second, it seems that due to the nature of ReLU-NNs on a scale finer than the given data resolution, structures in the image are promoted which are attractive to humans. It is not clear to us why this is the case and further studies in this direction might be needed.
In order to guarantee the existence of a (neural network) solution we introduced the constrained problem (9). In particular (9) has a solution in $\mathcal{H}$ as long as $c<\infty$. However, the constraint $|\theta|_{\infty}\leq c$ seems mainly theoretical. In particular, since our solution process relies on an iterative (stochastic) method, one could set $c=\infty$ in practice. Numerically we do not expect to reach the analytic solution exactly anyway, but only approximate it. In the case that the solution space is $\mathcal{H}$ we would obtain a solution in the space of piece-wise affine functions which is dense in the space $BV$. Hence, solving the optimization problem over $\mathcal{H}$ seems numerically sufficient, as the numerical solution in $\mathcal{H}$ can arbitrarily closely approximate the exact $BV$ solution. In fact a stopping criterion, as the number of iterations, will always allow the numerical process to terminate with an approximation.

### v2

@@ 5.1.3 Evaluation of the error estimate
for all $v\in L^{2}(\Omega)$, where $\rho_{1}(v):=2\|v-g\|_{L^{2}(\Omega)}$ and $\rho_{2}(v):=\frac{1}{\alpha_{2}}\|\xi_{G}\|_{L^{2}(\Omega)}$. In order to investigate the nature of the error estimate and how the two terms $\rho_{1}$ and $\rho_{2}$ behave, we depict $\rho$, $\rho_{1}$ and $\rho_{2}$ together with the real error in figure 8. Note that in figure 8 on the $x$-axes we have the number of updates of the approximation, i.e. the iterations at which the loss was reduced, and not the number of iterations. Exemplary we depict the error estimate for the example in section 5.1.1 with $c=100$ and for the example in section 5.1.2 with $c=10000$, using $\operatorname{TV}_{2}^{h}$ with Huber-regularization and $\operatorname{TV}_{2,1}^{h}$ with $\gamma$-lifting. A similar behavior can be observed for the other settings of sections 5.1.1 and 5.1.2. In figure 8 we see that while the error estimate indeed always holds, as expected, it is governed by $\rho_{2}$. In fact $\rho_{2}$ leads to an oscillatory behavior of the error estimate $\rho$, which does not seem to decrease although the real error is. This behavior does not really seem to make the error estimate practical, for example, to be used in a stopping criterion.
@@ 5.2 Applications
We consider three applications, namely denoising, inpainting and deblurring. In all the applications the image is scaled to the unit square $[0,1]^{2}$. In order to recover the original image from the corrupted observation we minimize (19). Thereby in all the experiments we choose a forward difference discretisation of the gradient operator with incorporated homogeneous Neumann boundary conditions. Further we choose $\lambda=1$ and $\operatorname{TV}_{2}^{h}$ with $\gamma$-lifting. That is the last term in (19), i.e. $\frac{\lambda}{N}\sum_{x\in\Omega^{h}}|\nabla^{h}u_{\theta}(x)|_{\blacktriangleright}$, is set to be $\frac{1}{N}\sum_{x\in\Omega^{h}}|\nabla_{\text{F}}^{h}u_{\theta}(x)|_{2,\gamma}$, where $\gamma=10^{-10}$. The solution space $\mathcal{H}_{c}^{M}$ consists of ReLU-NNs with 3 hidden layers, where each layer has 128 neurons, and we set $c=10000$. Adam with learning rate 0.001 is used to find an approximation and the training is always stopped after 300001 iterations. The parameters $\alpha_{1}$ and $\alpha_{2}$ are selected differently depending on the application and specified at the respective places. However, they are not optimized in any way and only chosen such that the effect of the total variation can be observed. Next, we describe the image reconstruction tasks before we present and discuss the respective results.
@@ Results
The obtained restoration for all applications on the grid of the observed image can be seen in the second column of figures 11 and 12. Since a neural network is a continuous function not restricted to a particular grid, the obtained solutions can be plotted on any discretisation of $\Omega$. We plot the obtained solutions on the grid of the observed image, i.e. on a grid with $129\times 129$ pixels for the disk-image and $256\times 256$ pixels for the building-image, see second column of figures 11 and 12, and on a 10 times finer grid, i.e. on a grid with $1290\times 1290$ pixels for the disk-image and $2560\times 2560$ pixels for the building-image, see fourth column of figures 11 and 12. In order to see the difference in representation of the solutions we zoom-in in one part of the images, see third and fifth column of figures 11 and 12. Here we observe that the solutions on a fine grid seems to be well adjusted to the nature of the original signal to be restored, e.g. the curvature of the disk seems to be preserved. However, we emphasize that according to propositions 4.7 and 4.8 on a fixed grid the solution of (19) is the same as the respective finite difference solution, if it exists, in the sense that the values of the solutions on the grid points are the same. Hence the effects of the neural network’s design are only apparent or relevant at a level of detail finer than what is captured by the current data resolution.
@@ 6 Conclusion and discussion
We have seen that the considered neural network approach changes the nature of the original problem. In particular, while the $L^{1}$-$L^{2}$-TV model is convex, its neural network version (4) is not convex anymore. Further we observe that even if the original problem has a solution the resulting neural network optimization problem does in general not, cf. example 3.3. Hence, it seems that the neural network approach renders the problem more complicated to deal with. Note that the non-existence of a solution can be easily resolved, as we did by introducing (9). However, the non-convexity seems to be persistent and we are not aware how to circumvent this property. Nevertheless, this additional complexity seems to come with a benefit. First, while optimizing on a rather coarse grid, we actually obtain a solution in a continuum, i.e. the solution is mesh-free and continuous, allowing us to depict the solution on an arbitrary fine grid and any domain. Second, it seems that due to the nature of ReLU-NNs on a scale finer than the given data resolution, structures in the image are promoted which are attractive to humans. It is not clear to us why this is the case and further studies in this direction might be needed.
In order to guarantee the existence of a (neural network) solution we introduced the constrained problem (9). In particular (9) has a solution in $\mathcal{H}^{M}$ as long as $c<\infty$. However, the constraint $|\theta|_{\infty}\leq c$ seems mainly theoretical. In particular, since our solution process relies on an iterative (stochastic) method, numerically we do not expect to reach the analytic solution exactly anyway, but only approximate it. In the case that the solution space is $\mathcal{H}^{M}$, i.e. $c=\infty$, we would obtain an approximation in the space of piecewise affine functions which is dense in the space $BV$. Hence, solving the optimization problem over $\mathcal{H}^{M}$ seems numerically sufficient, as the numerical solution in $\mathcal{H}^{M}$ can arbitrarily closely approximate the exact $BV$ solution, assuming that $M$ is sufficiently large, and a stopping criterion, as the number of iterations, will always allow the numerical process to terminate with an approximation.